OpenStudy (anonymous):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Make 6 squares. The squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing the tile numbered 5 from the first bag and an odd tile from the second bag?

Vocaloid (vocaloid):

let's break the problem down into smaller pieces 1. what is the probability of drawing a tile numbered 5 from the first bag?

OpenStudy (anonymous):

1/6

Vocaloid (vocaloid):

right! now let's move to the next part 2. what is the probability of drawing an odd tile from the bag?

OpenStudy (anonymous):

1/2

Vocaloid (vocaloid):

right! now multiply the two probabilities together (1/6)*(1/2) = ?

OpenStudy (anonymous):

3/6

Vocaloid (vocaloid):

uh, not quite, we need to multiply them together 2*6 = ?

OpenStudy (anonymous):

12/36

OpenStudy (anonymous):

im so confused

Vocaloid (vocaloid):

not quite, all you need to do is multiply the numerators, then multiply the denominators hint: (A/B)*(C/D) = A*C/(B*D)

Vocaloid (vocaloid):



OpenStudy (anonymous):

3/36